OpenStudy (bossimbacon):

How many solutions does the equation have? 4x+1=5+2(2x-4) A. one solution B. infinite solutions C. no solution

OpenStudy (anonymous):

4x + 1 = 5 + 4x - 8 4x +1 = 4x - 3 and you can guess the option

OpenStudy (bossimbacon):

ummm one solution?